Consider the reaction: A(g)+12B(g)→2C(g).

When C is increasing at a rate of 4.0×10−2 M⋅s−1, how fast is B decreasing?
Express your answer using two significant figures.
Δ[B]/Δt = ? M⋅s−1
1
Expert's answer
2020-06-16T15:04:01-0400

-1/12∆B/∆t=1/2∆C/∆t

=-6(4.0*10^-2)

=24*10-2 M⋅s−1